If net sales revenue and the average book value of fixed assets both rise 5%:

Multiple Choice

A. the fixed asset turnover ratio will rise.

B. the fixed asset turnover ratio will fall.

C. the fixed asset turnover ratio will stay the same.

D. the impact on the fixed asset turnover ratio cannot be determined since the beginning values are unknown.
Answer #1
The fixed asset turnover ratio will stay the same.
Option C is correct
Workings:
Let's assume the current net sales revenue to be $200000 and average book value of fixed assets to be $100000
Fixed asset turnover ratio = 200000/100000 = 2.00
After 5% increase, Net sales revenue = 200000*1.05 = $210000; Fixed asset turnover ratio = 100000*1.05 = $105000
Fixed asset turnover ratio = 210000/105000 = 2.00
The fixed asset turnover ratio is same before and after increase.